OpenStudy (anonymous):

Please help me! Will give medal :) Gino designed a garden for his rectangular backyard. The design he drew on paper is 7 inches wide. The width of his backyard is forty-two feet. If his drawing is to scale and is 8 inches long, how many feet long is his backyard? Write a proportion using words. Then, use the proportion to calculate the yard length. Show your work.

OpenStudy (anonymous):

@proftom

OpenStudy (anonymous):

Seven inches is to forty two feet as eight inches is two the length of his backyard in feet. 7/42=8/x 336=7x x=48 feet 48/3= 16 yards (long)

OpenStudy (anonymous):

thank you so much!... is that all ?

OpenStudy (anonymous):

if the backyard is 42 ft and in the drawing, this is 7 in, then each inch on the drawing is 6 ft. (42ft/7in = 6 ft/in) So an 8in long yard would be 8in * 6f ft/in = 48 ft.

OpenStudy (anonymous):

um... thank u too.. but which answer should i choose yours or slyougdale's

OpenStudy (merchandize):

Gino designed a garden for his rectangular backyard. The design he drew on paper is 7 inches wide. The width of his backyard is 42 feet. 7 : 42 x 12 or 1: 72 If his drawing is to scale and is 8 inches long, how many feet long is his backyard? His backyard is 8 inches long on the drawing. 8 inches represent 8 times 72 inches on the ground Hence the yard length is 576 inches or 48 feet.

OpenStudy (anonymous):

thank u too... so your answer and Harsh123 are right and not slyoungdale

OpenStudy (anonymous):

@proftom

OpenStudy (anonymous):

Sly is right. the yard is 48 feet long, but the question asks for the answer to be in yards. 48 ft = 16 yards

OpenStudy (anonymous):

to be safe, i would list both answers: the length of the yard is 48 feet or 16 yards.

OpenStudy (anonymous):

ok thank you!!